# Minimizing Human Intervention in Online Classification

### v1

@@ F.1.2 Evaluation of CC and Comparison with GHC
where $\sigma=0.1$ and $C=108$. As discussed in Appendix E.5, this rule is best suited for the Euclidean setting $\mathcal{E}=\mathcal{I}^{d}$. We therefore treat $C$ as a tunable parameter to find a practical balance in all settings. The results for various values of CC are reported in Tables 3 and 4. To allow a direct comparison with the performance of previous algorithms, we only report in “C” the number of expert calls after each label has been observed at least once (this corresponds to the number of expert calls in the second phase for the GHC and $\operatorname{SKM}$ algorithms).
@@
Dim Dist. $\tau=0.00$ $\tau=0.20$ $\tau=0.40$ $\tau=0.60$ $\tau=0.80$ $\tau=0.95$ $\tau=1.00$ 1 Unif. C: 0± $0\,\pm\,$0 M: $375\,\pm\,126$ R: $4160\,\pm\,1384$ C: $124\,\pm\,3$ M: 2± $2\,\pm\,$2 R: $302\,\pm\,19$ C: $59\,\pm\,3$ M: 3± $3\,\pm\,$3 R: $186\,\pm\,33$ C: $35\,\pm\,3$ M: $12\,\pm\,3$ R: $237\,\pm\,33$ C: $21\,\pm\,4$ M: $34\,\pm\,10$ R: $445\,\pm\,110$ C: $12\,\pm\,2$ M: $101\,\pm\,33$ R: $1168\,\pm\,363$ C: 0± $0\,\pm\,$0 M: $375\,\pm\,126$ R: $4160\,\pm\,1384$ Mix. C: 0± $0\,\pm\,$0 M: $823\,\pm\,389$ R: $9071\,\pm\,4275$ C: $136\,\pm\,9$ M: 1± $1\,\pm\,$1 R: $307\,\pm\,21$ C: $65\,\pm\,5$ M: $9\,\pm\,5$ R: $245\,\pm\,50$ C: $40\,\pm\,1$ M: $21\,\pm\,7$ R: $333\,\pm\,78$ C: $27\,\pm\,1$ M: $42\,\pm\,14$ R: $532\,\pm\,151$ C: $15\,\pm\,2$ M: $155\,\pm\,31$ R: $1753\,\pm\,337$ C: 0± $0\,\pm\,$0 M: $823\,\pm\,389$ R: $9071\,\pm\,4275$ 4 Unif. C: 0± $0\,\pm\,$0 M: $2155\,\pm\,181$ R: $23729\,\pm\,1988$ C: $4723\,\pm\,8$ M: 0± $0\,\pm\,$0 R: $9477\,\pm\,17$ C: $3530\,\pm\,16$ M: $12\,\pm\,3$ R: $7224\,\pm\,65$ C: $2190\,\pm\,24$ M: $90\,\pm\,6$ R: $5396\,\pm\,95$ C: $1068\,\pm\,29$ M: $348\,\pm\,16$ R: $5988\,\pm\,188$ C: $319\,\pm\,16$ M: $894\,\pm\,30$ R: $10496\,\pm\,305$ C: 0± $0\,\pm\,$0 M: $2155\,\pm\,181$ R: $23729\,\pm\,1988$ Mix. C: 0± $0\,\pm\,$0 M: $422\,\pm\,199$ R: $4666\,\pm\,2188$ C: $3450\,\pm\,45$ M: 0± $0\,\pm\,$0 R: $6928\,\pm\,81$ C: $1575\,\pm\,31$ M: $2\,\pm\,1$ R: $3196\,\pm\,49$ C: $760\,\pm\,26$ M: $13\,\pm\,4$ R: $1690\,\pm\,79$ C: $343\,\pm\,22$ M: $59\,\pm\,7$ R: $1366\,\pm\,99$ C: $98\,\pm\,9$ M: $187\,\pm\,24$ R: $2283\,\pm\,267$ C: 0± $0\,\pm\,$0 M: $422\,\pm\,199$ R: $4666\,\pm\,2188$ 10 Unif. C: 0± $0\,\pm\,$0 M: $2731\,\pm\,236$ R: $30075\,\pm\,2577$ C: $4983\,\pm\,10$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4972\,\pm\,11$ M: 0± $0\,\pm\,$0 R: $9978\,\pm\,2$ C: $4704\,\pm\,24$ M: $18\,\pm\,4$ R: $9641\,\pm\,38$ C: $3235\,\pm\,22$ M: $276\,\pm\,7$ R: $9544\,\pm\,91$ C: $1027\,\pm\,29$ M: $1242\,\pm\,24$ R: $15751\,\pm\,316$ C: 0± $0\,\pm\,$0 M: $2731\,\pm\,236$ R: $30075\,\pm\,2577$ Mix. C: 0± $0\,\pm\,$0 M: 1± $1\,\pm\,$1 R: $40\,\pm\,12$ C: $4635\,\pm\,21$ M: 0± $0\,\pm\,$0 R: $9294\,\pm\,45$ C: $1081\,\pm\,24$ M: 0± $0\,\pm\,$0 R: $2187\,\pm\,54$ C: $174\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $372\,\pm\,13$ C: $26\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $78\,\pm\,15$ C: 1± $1\,\pm\,$1 M: 0± $0\,\pm\,$0 R: $29\,\pm\,10$ C: 0± $0\,\pm\,$0 M: 1± $1\,\pm\,$1 R: $40\,\pm\,12$ 50 Unif. C: 0± $0\,\pm\,$0 M: $3344\,\pm\,52$ R: $36828\,\pm\,570$ C: $4978\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4978\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4978\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4974\,\pm\,4$ M: $2\,\pm\,1$ R: $10019\,\pm\,8$ C: $3551\,\pm\,32$ M: $738\,\pm\,25$ R: $15268\,\pm\,221$ C: 0± $0\,\pm\,$0 M: $3344\,\pm\,52$ R: $36828\,\pm\,570$ Mix. C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $24\,\pm\,9$ C: $4987\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $9998\,\pm\,3$ C: $254\,\pm\,17$ M: 0± $0\,\pm\,$0 R: $532\,\pm\,31$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $24\,\pm\,9$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $24\,\pm\,9$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $24\,\pm\,9$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $24\,\pm\,9$
@@
Dim Dist. $\tau=0.00$ $\tau=0.20$ $\tau=0.40$ $\tau=0.60$ $\tau=0.80$ $\tau=1.00$ 2 Unif. C: 0± $0\,\pm\,$0 M: $315\,\pm\,148$ R: $3524\,\pm\,1630$ C: $142\,\pm\,7$ M: $2\,\pm\,1$ R: $371\,\pm\,15$ C: $69\,\pm\,5$ M: $5\,\pm\,3$ R: $254\,\pm\,42$ C: $39\,\pm\,3$ M: $18\,\pm\,6$ R: $340\,\pm\,73$ C: $24\,\pm\,2$ M: 44± $4\,\pm\,11$ R: $596\,\pm\,144$ C: 0± $0\,\pm\,$0 M: $315\,\pm\,148$ R: $3524\,\pm\,1630$ Mix. C: 0± $0\,\pm\,$0 M: $345\,\pm\,77$ R: $3812\,\pm\,849$ C: $100\,\pm\,5$ M: 1± $1\,\pm\,$1 R: $231\,\pm\,17$ C: $50\,\pm\,5$ M: $3\,\pm\,1$ R: $155\,\pm\,19$ C: $30\,\pm\,2$ M: $9\,\pm\,3$ R: $180\,\pm\,30$ C: $18\,\pm\,3$ M: 22± $2\,\pm\,9$ R: $302\,\pm\,100$ C: 0± $0\,\pm\,$0 M: $345\,\pm\,77$ R: $3812\,\pm\,849$ 4 Unif. C: 0± $0\,\pm\,$0 M: $1939\,\pm\,308$ R: $21348\,\pm\,3381$ C: $4117\,\pm\,19$ M: 1± $1\,\pm\,$1 R: $8272\,\pm\,46$ C: $2526\,\pm\,21$ M: 22± $2\,\pm\,6$ R: $5319\,\pm\,79$ C: $1425\,\pm\,9$ M: $103\,\pm\,11$ R: $4005\,\pm\,126$ C: 696± $96\,\pm\,$9 M: $307\,\pm\,9$ R: $4793\,\pm\,92$ C: 0± $0\,\pm\,$0 M: $1939\,\pm\,308$ R: $21348\,\pm\,3381$ Mix. C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,7$ C: $121\,\pm\,10$ M: 0± $0\,\pm\,$0 R: $264\,\pm\,19$ C: $20\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $62\,\pm\,11$ C: $4\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $31\,\pm\,3$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,6$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,7$ 10 Unif. C: 0± $0\,\pm\,$0 M: $3039\,\pm\,213$ R: $33446\,\pm\,2337$ C: $4990\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4983\,\pm\,5$ M: $1\,\pm\,0$ R: $9993\,\pm\,7$ C: $4766\,\pm\,7$ M: $17\,\pm\,5$ R: $9743\,\pm\,51$ C: $3412\,\pm\,25$ M: $284\,\pm\,12$ R: $9971\,\pm\,142$ C: 0± $0\,\pm\,$0 M: $3039\,\pm\,213$ R: $33446\,\pm\,2337$ Mix. C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,9$ C: $1769\,\pm\,45$ M: 0± $0\,\pm\,$0 R: $3560\,\pm\,86$ C: $105\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $231\,\pm\,7$ C: $3\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $28\,\pm\,8$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,9$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,9$ 50 Unif. C: 0± $0\,\pm\,$0 M: $3549\,\pm\,83$ R: $39061\,\pm\,911$ C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4987\,\pm\,2$ M: 1± $1\,\pm\,$1 R: $10005\,\pm\,10$ C: 0± $0\,\pm\,$0 M: $3549\,\pm\,83$ R: $39061\,\pm\,911$ Mix. C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $26\,\pm\,12$ C: $4987\,\pm\,6$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4983\,\pm\,7$ M: 0± $0\,\pm\,$0 R: $9992\,\pm\,1$ C: $1287\,\pm\,44$ M: 0± $0\,\pm\,$0 R: $2600\,\pm\,97$ C: $12\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $49\,\pm\,10$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $26\,\pm\,12$
Table 7: Voronoi regret of all algorithms for each experimental setup ($T=5000$). Results for CC and GHC are reported for the best-performing threshold for that setup.

### v2

@@ F.1.2 Evaluation of CC and Comparison with GHC
where $\sigma=0.1$ and $C=108$. As discussed in Appendix E.5, this rule is best suited for the Euclidean setting $\mathcal{E}=\mathcal{I}^{d}$. We therefore treat $C$ as a tunable parameter to find a practical balance in all settings. The results for various values of CC are reported in Tables 3 and 4. To allow a direct comparison with the performance of previous algorithms, we only report in “C” the number of expert calls after each label has been observed at least once (this corresponds to the number of expert calls in the second phase for the GHC and SKM algorithms).
@@
Dim Dist. $\tau=0.00$ $\tau=0.20$ $\tau=0.40$ $\tau=0.60$ $\tau=0.80$ $\tau=0.95$ $\tau=1.00$ 1 Unif. C: $4983\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $9996\,\pm\,3$ C: $124\,\pm\,3$ M: 2± $2\,\pm\,$2 R: $302\,\pm\,19$ C: $59\,\pm\,3$ M: 3± $3\,\pm\,$3 R: $186\,\pm\,33$ C: $35\,\pm\,3$ M: $12\,\pm\,3$ R: $237\,\pm\,33$ C: $21\,\pm\,4$ M: $34\,\pm\,10$ R: $445\,\pm\,110$ C: $12\,\pm\,2$ M: $101\,\pm\,33$ R: $1168\,\pm\,363$ C: 0± $0\,\pm\,$0 M: $375\,\pm\,126$ R: $4160\,\pm\,1384$ Mix. C: $4987\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $9994\,\pm\,5$ C: $136\,\pm\,9$ M: 1± $1\,\pm\,$1 R: $307\,\pm\,21$ C: $65\,\pm\,5$ M: $9\,\pm\,5$ R: $245\,\pm\,50$ C: $40\,\pm\,1$ M: $21\,\pm\,7$ R: $333\,\pm\,78$ C: $27\,\pm\,1$ M: $42\,\pm\,14$ R: $532\,\pm\,151$ C: $15\,\pm\,2$ M: $155\,\pm\,31$ R: $1753\,\pm\,337$ C: 0± $0\,\pm\,$0 M: $823\,\pm\,389$ R: $9071\,\pm\,4275$ 4 Unif. C: $4986\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4723\,\pm\,8$ M: 0± $0\,\pm\,$0 R: $9477\,\pm\,17$ C: $3530\,\pm\,16$ M: $12\,\pm\,3$ R: $7224\,\pm\,65$ C: $2190\,\pm\,24$ M: $90\,\pm\,6$ R: $5396\,\pm\,95$ C: $1068\,\pm\,29$ M: $348\,\pm\,16$ R: $5988\,\pm\,188$ C: $319\,\pm\,16$ M: $894\,\pm\,30$ R: $10496\,\pm\,305$ C: 0± $0\,\pm\,$0 M: $2155\,\pm\,181$ R: $23729\,\pm\,1988$ Mix. C: $4986\,\pm\,7$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $3450\,\pm\,45$ M: 0± $0\,\pm\,$0 R: $6928\,\pm\,81$ C: $1575\,\pm\,31$ M: $2\,\pm\,1$ R: $3196\,\pm\,49$ C: $760\,\pm\,26$ M: $13\,\pm\,4$ R: $1690\,\pm\,79$ C: $343\,\pm\,22$ M: $59\,\pm\,7$ R: $1366\,\pm\,99$ C: $98\,\pm\,9$ M: $187\,\pm\,24$ R: $2283\,\pm\,267$ C: 0± $0\,\pm\,$0 M: $422\,\pm\,199$ R: $4666\,\pm\,2188$ 10 Unif. C: $4983\,\pm\,10$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4983\,\pm\,10$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4972\,\pm\,11$ M: 0± $0\,\pm\,$0 R: $9978\,\pm\,2$ C: $4704\,\pm\,24$ M: $18\,\pm\,4$ R: $9641\,\pm\,38$ C: $3235\,\pm\,22$ M: $276\,\pm\,7$ R: $9544\,\pm\,91$ C: $1027\,\pm\,29$ M: $1242\,\pm\,24$ R: $15751\,\pm\,316$ C: 0± $0\,\pm\,$0 M: $2731\,\pm\,236$ R: $30075\,\pm\,2577$ Mix. C: $4988\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4635\,\pm\,21$ M: 0± $0\,\pm\,$0 R: $9294\,\pm\,45$ C: $1081\,\pm\,24$ M: 0± $0\,\pm\,$0 R: $2187\,\pm\,54$ C: $174\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $372\,\pm\,13$ C: $26\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $78\,\pm\,15$ C: 1± $1\,\pm\,$1 M: 0± $0\,\pm\,$0 R: $29\,\pm\,10$ C: 0± $0\,\pm\,$0 M: 1± $1\,\pm\,$1 R: $40\,\pm\,12$ 50 Unif. C: $4978\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4978\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4978\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4978\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4974\,\pm\,4$ M: $2\,\pm\,1$ R: $10019\,\pm\,8$ C: $3551\,\pm\,32$ M: $738\,\pm\,25$ R: $15268\,\pm\,221$ C: 0± $0\,\pm\,$0 M: $3344\,\pm\,52$ R: $36828\,\pm\,570$ Mix. C: $4988\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4987\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $9998\,\pm\,3$ C: $254\,\pm\,17$ M: 0± $0\,\pm\,$0 R: $532\,\pm\,31$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $24\,\pm\,9$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $24\,\pm\,9$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $24\,\pm\,9$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $24\,\pm\,9$
@@
Dim Dist. $\tau=0.00$ $\tau=0.20$ $\tau=0.40$ $\tau=0.60$ $\tau=0.80$ $\tau=1.00$ 2 Unif. C: $4969\,\pm\,10$ M: 0± $0\,\pm\,$0 R: $9998\,\pm\,1$ C: $142\,\pm\,7$ M: $2\,\pm\,1$ R: $371\,\pm\,15$ C: $69\,\pm\,5$ M: $5\,\pm\,3$ R: $254\,\pm\,42$ C: $39\,\pm\,3$ M: $18\,\pm\,6$ R: $340\,\pm\,73$ C: $24\,\pm\,2$ M: 44± $4\,\pm\,11$ R: $596\,\pm\,144$ C: 0± $0\,\pm\,$0 M: $315\,\pm\,148$ R: $3524\,\pm\,1630$ Mix. C: $4988\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $9996\,\pm\,3$ C: $100\,\pm\,5$ M: 1± $1\,\pm\,$1 R: $231\,\pm\,17$ C: $50\,\pm\,5$ M: $3\,\pm\,1$ R: $155\,\pm\,19$ C: $30\,\pm\,2$ M: $9\,\pm\,3$ R: $180\,\pm\,30$ C: $18\,\pm\,3$ M: 22± $2\,\pm\,9$ R: $302\,\pm\,100$ C: 0± $0\,\pm\,$0 M: $345\,\pm\,77$ R: $3812\,\pm\,849$ 4 Unif. C: $4989\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4117\,\pm\,19$ M: 1± $1\,\pm\,$1 R: $8272\,\pm\,46$ C: $2526\,\pm\,21$ M: 22± $2\,\pm\,6$ R: $5319\,\pm\,79$ C: $1425\,\pm\,9$ M: $103\,\pm\,11$ R: $4005\,\pm\,126$ C: 696± $96\,\pm\,$9 M: $307\,\pm\,9$ R: $4793\,\pm\,92$ C: 0± $0\,\pm\,$0 M: $1939\,\pm\,308$ R: $21348\,\pm\,3381$ Mix. C: $4989\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $121\,\pm\,10$ M: 0± $0\,\pm\,$0 R: $264\,\pm\,19$ C: $20\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $62\,\pm\,11$ C: $4\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $31\,\pm\,3$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,6$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,7$ 10 Unif. C: $4990\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4990\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4983\,\pm\,5$ M: $1\,\pm\,0$ R: $9993\,\pm\,7$ C: $4766\,\pm\,7$ M: $17\,\pm\,5$ R: $9743\,\pm\,51$ C: $3412\,\pm\,25$ M: $284\,\pm\,12$ R: $9971\,\pm\,142$ C: 0± $0\,\pm\,$0 M: $3039\,\pm\,213$ R: $33446\,\pm\,2337$ Mix. C: $4989\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $1769\,\pm\,45$ M: 0± $0\,\pm\,$0 R: $3560\,\pm\,86$ C: $105\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $231\,\pm\,7$ C: $3\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $28\,\pm\,8$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,9$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,9$ 50 Unif. C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4987\,\pm\,2$ M: 1± $1\,\pm\,$1 R: $10005\,\pm\,10$ C: 0± $0\,\pm\,$0 M: $3549\,\pm\,83$ R: $39061\,\pm\,911$ Mix. C: $4987\,\pm\,6$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4987\,\pm\,6$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4983\,\pm\,7$ M: 0± $0\,\pm\,$0 R: $9992\,\pm\,1$ C: $1287\,\pm\,44$ M: 0± $0\,\pm\,$0 R: $2600\,\pm\,97$ C: $12\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $49\,\pm\,10$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $26\,\pm\,12$
Table 7: Voronoi regret of all algorithms for each experimental setup ($T=5000$). Results for $C$, GHC and AMP are reported for the best-performing threshold for that setup.